The difference between two numbers is 13. Two times the larger is 17 less than 3 times the smaller. What are the numbers?
Answer by drk(1908)

The difference between two numbers is 13 translates to
X - Y = 13
Two times the larger is 17 less than 3 times the smaller translates to
2X = 3Y - 17.
We can solve by substitution and get
2(y+13) = 3y - 17.
2y + 26 = 3y - 17
43 = y
x = 56
(56, 43)